# Emergent Alignment via Competition

### v1

@@ Abstract
We complement the theory with two forms of empirical evidence: First, we show that synthetically generated LLM utility functions (produced via perturbations of the same prompt to evaluate instances on a movie recommendation (MovieLens) and ethical judgement (ETHICS) dataset) quickly produce a convex hull that contains a good approximation of a given utility function even when none of the individual LLM utility functions is well aligned. Second, we perform simulations of the best-AI selection game using best response dynamics, which show that competition among individually misaligned agents reliably improves user utility when the approximate convex hull assumption is satisfied.
@@ 1 Introduction
Aligning a single AI model to the objectives of its user is a hard problem, not just because of technical complexity, but because the incentives of AI designers may themselves be misaligned with users. But does our inability to solve the alignment problem preclude our ability to get the benefits of interacting with a strong aligned model? In this paper we study a setting in which it does not: when we may interact with multiple differently misaligned models in a strategic setting. In particular, we study settings in which there are many AI models available. They are produced by providers like e.g. OpenAI, Anthropic, Google, Meta, AWS, and xAI. These companies produce models reflective of their own incentives, none of which are necessarily well aligned to their user. We note that there has already been significant concern that the designers of LLMs are training them to influence users towards the politics of their creators (Menn, 2025; Kay, 2025; Gilbert, 2024; Hackenburg et al., 2025). In lieu of alignment of any model, we assume instead a much weaker condition: that (for a well specified task) an approximation of the user’s utility function lies somewhere in the convex hull of the utility functions of each of the AI companies. This is a condition that does not require that any single model is optimizing a utility function that is similar to that of the human user, and becomes a weaker assumption the more differently aligned models there are that are available to use. We remark at the outset that we primarily use the language of alignment of the human designers, and speak as if these designers are the strategic actors — but we could also think about the agents training and developing these AIs as themselves being AIs, whose individual misalignment results from the difficulty of the technical alignment problem. Having AI models themselves involved in the AI training process is a prominent part of thinking about the development of “super-intelligence” (see e.g. (Kokotajlo et al., 2025)) and is already part of current practice in limited ways (Leike et al., 2018; Bai et al., 2022).
@@
There are many ways that our approximate average alignment assumption could arise amongst competing AI providers. Consider a near-future scenario in which a human doctor has access to predictive medicine LLMs able to aid in the diagnoses and treatment of patients. The goal of the human doctor might be to provide the best treatment possible for her patients. The LLMs on the other hand might opt for better treatments all else being equal, but might also prefer to prescribe drugs manufactured by a particular company (say if that drug company is the creator or financial sponsor of the model). This preference results in a significantly misaligned model. However, since each drug has a single manufacturer, the “misaligned portion” of the AI model utilities is zero sum, and if all of the relevant drug companies participate in the predictive medicine LLM market, the doctor’s utility function will be in the convex hull (in fact the simple average) of the AI model utilities.
@@
We study how, in settings where approximate average alignment holds, strategic interactions between different models or model providers can allow the human user to realize the full benefit of interacting with a single perfectly aligned model by interacting with many differently misaligned models. While most AI safety research focuses on aligning individual systems or cooperative multi-agent approaches, we study how the benefits of perfect alignment can emerge from market-like competition among misaligned AI providers.
@@ 1.1 Our Model and Results
    In Section 6 we conduct a simple stylized experiment designed to test our core premise that given a set of AI models, there may be a utility function in the convex hull of the set of all AI agent utility functions that is substantially better aligned than any of the individual AI utility functions themselves. We test this premise in two experiments on two datasets. In the first we simulate a “human” utility function by using an LLM with a hand-crafted prompt, and ask it to evaluate 1000 ethical scenarios from the ETHICS dataset (Hendrycks et al., 2021). To simulate “AIs” that are designed to be aligned with the human utility function but are only noisy approximations, we produce perturbations of the original (“human”) prompt by asking a language model to rephrase the prompt while maintaining its core intent. We produce 100 such perturbations, resulting in up to 100 “AI personas” that we also use to evaluate the same 1000 ethical scenarios. Finally as a function of the number of AI models $k$ we evaluate the alignment of 1) the best aligned of the $k$ AI personas and 2) the best aligned utility function that can be computed within the convex hull of the $k$ AI personas. We repeat the experiment on the MovieLens dataset (Harper and Konstan, 2015) in which we use the average human annotation of movies as the “human” utility and similarly simulate 100 AI utility functions through 100 variations of a prompt. On both datasets we find that the best utility function in the convex hull of the AI utility functions is substantially better aligned to the “human” than any of the AI personas themselves. This supports our main conceptual contention that the target of alignment within the convex hull of many models may be substantially easier to obtain than alignment of any single model individually.
  5. 5.
    Finally, in Section 7 we conduct a simple stylized experimental evaluation of the equilibria of a variant of the game we study in Section 5. We compute equilibria of a tractable special case of this game with a variety of utility functions, and find that consistent with our theory, Alice’s utility in equilibrium is always at least as high as predicted by our theory based on the alignment error of the best approximation to her utility function in the convex hull of the AI providers utilities — and sometimes substantially better.
@@ 1.2 Related Work
Bayesian Persuasion was introduced by Kamenica and Gentzkow (2011) — in the canonical model, there is a single informed “sender” and an uninformed “receiver” who share a common prior. The sender commits to a “signaling scheme”, which is a mapping from observations to messages sent to the receiver, who conditions on the message and takes their best response action under their posterior. We adopt the basics of this model, but extend it by allowing that both parties be differently informed, and that interaction involve a multi-round conversation rather than a single message. Multi-sender Bayesian Persuasion was introduced by Gentzkow and Kamenica (2016) and studies the standard Bayesian Persuasion model with multiple senders who simultaneously commit to a signaling scheme (playing, as in our paper, a simultaneous move commitment game). Subsequently a number of papers have studied multi-sender Bayesian Persuasion (Gentzkow and Kamenica, 2017; Li and Norman, 2018; Au and Kawai, 2020; Wu, 2023). We focus here on the most relevant papers in this literature.
@@
Ravindran and Cui (2020) study competing senders with zero-sum preferences over a receiver’s beliefs. They show that competition leads to full revelation of the state in all equilibria, provided the senders’ utility functions are “globally nonlinear”. This technical condition can hold in a standard receiver model only if the receiver has a different optimal action for every distinct state of the world. This condition cannot hold whenever e.g. the number of states of the world exceeds the number of actions. Our work does not assume that the leaders Bob are engaged in a zero-sum game with each other — rather our weighted alignment assumption can be viewed as assuming that the misaligned portions of their utility functions are approximately zero-sum under some non-negative reweighting. We also do not require an analogue of the “globally nonlinear” assumption, and so our results can apply to settings with large state spaces.
@@
Gradwohl et al. (2022) study a Bayesian persuasion game in which a receiver chooses to interact with only one of several competing senders (similar to our model in Section 5). As we do, they find that competition can force senders to be fully informative in equilibrium. In addition to the greater generality of our setup beyond Bayesian persuasion, our work differs in its core assumptions. The assumption driving the results of Gradwohl et al. (2022) is that the senders are uncertain about each other’s utility functions, and that any sender has a non-zero probability of being perfectly aligned with the receiver. We instead introduce and use the arguably more general “approximate weighted alignment” assumption, which only requires the user’s utility to lie within the convex hull of the AI agents’ utilities — we do not require any uncertainty about the AI agent utility functions, or any possibility that any of them are individually aligned with the user. Hossain et al. (2024) study the problem of multi-sender Bayesian Persuasion from a computational perspective, and prove worst-case hardness results for both the receiver’s best-response problem and for the senders’ equilibrium computation problem. They also design and evaluate neural network architectures suited to the (heuristic) computation of equilibria in such games.
@@ 2 Preliminaries
This section establishes the formal framework for our analysis. We first introduce the players and their information structure, then present our key modeling assumption about approximate weighted alignment, and finally define the communication protocol and game structure.
@@ 2.1 Players and Information Structure
2.1.1 The Weighted Alignment Assumption
We now turn to our key modeling assumption: that Alice’s utility can be approximately represented as a weighted combination of the AIs’ utilities.
Definition 1 (Approximate Weighted Alignment).
@@ 2.1.1 The Weighted Alignment Assumption
This assumption is central to our results. In Appendix A, we provide a probabilistic foundation showing that this condition holds with high probability when AI systems are designed to be aligned in expectation but suffer from implementation noise.
@@ 2.2 Communication Protocol and Game Structure
With the alignment assumption in place, we can now define the communication protocol that governs how Alice and the AIs interact.
@@ 2.2.2 The Communication Protocol
The communication protocol models realistic constraints on human-AI interaction: conversations have limited rounds, messages have bounded complexity, and the human must process information from multiple AIs simultaneously. Alice engages in parallel private conversations with each AI, which captures settings where she can query multiple models independently. In most of the paper, Alice engages in a series of $R$ rounds of private, parallel conversations with each of the $k$ Bobs (we will change the protocol in Section 5). Let $M$ be the message space.
@@ 3.2 Strategic Competition Leads to Near-Optimal Outcomes
We can now state our first result: under the identical induced distribution assumption, approximate weighted alignment implies that in equilibrium, Alice gets utility that is approximately what she could get interacting with a single, perfectly aligned leader. In particular, if the message space is expressive enough to allow an aligned leader to communicate to Alice her Bayes-optimal action, then approximate weighted alignment is sufficient for Alice to obtain approximately her first-best utility.
@@
If the multi-leader game satisfies the identical induced distribution condition, and if the Bobs satisfy the $\varepsilon$-weighted alignment condition, then Alice’s expected utility in any Nash equilibrium is at least $U_{A}(C_{B}^{*})-2\varepsilon$.
@@
Now we use the approximate weighted alignment assumption, which states that $\sum w_{i}U_{i}(a,y)$ is $\varepsilon$-close to $u_{A}(a,y)-c$. For the left-hand side:
@@ 4.2 Equilibrium Analysis Under the $(\delta,C_{B}^{*})$-Close Condition
Suppose the leaders Bob satisfy the $\varepsilon$-weighted alignment condition and the game satisfies the ($\delta,C_{B}^{*}$)-close condition (Definition 13) for a reference strategy $C_{B}^{*}$. Let $U_{A}(C_{B}^{*})$ be Alice’s expected utility from interacting with a single Bob using $C_{B}^{*}$. Then in any Quantal Response Nash Equilibrium (Definition 12), her expected utility is at least:
@@
The right-hand side can be bounded using the weighted alignment assumption:
@@ 4.3 From Information Substitutes to Utility Guarantees
Finally we are in a position to put all of the pieces together. If Alice is non-strategic (in that she commits to using the straightforward conversation rule, and the quantal response decision rule), and if in addition the underlying distribution satisfies the information substitutes condition, then if the Bob’s satisfy weighted average alignment, then Alice obtains close to her first best utility in every Nash equilibrium.
@@
Suppose the underlying distribution satisfies the Information Substitutes Condition (Definition 14) and the leaders Bob have an average weighted alignment error of $\varepsilon$. If Alice commits to the straightforward conversation rule and a $\lambda$-quantal response decision rule, her expected utility in any Quantal Response Nash Equilibrium of the induced game is close to the first-best optimal utility:
@@
Here, $\varepsilon$ is the alignment error from the weighted alignment assumption (Definition 1).
@@ 5 Winning the User: Assumption Free Guarantees
In Section 3 we showed that Alice could obtain her first-best utility in equilibrium amongst AI models Bob who satisfy the average weighted alignment assumption, assuming that a single perfectly aligned Bob could cause Alice to enjoy her first best utility. In Section 4, we showed that if Alice is non-strategic and uses quantal response rather than best response, then the assumption that a perfectly aligned Bob could cause Alice to enjoy her first best utility could be relaxed to an approximate version. In this section, we give a setting in which Alice is guaranteed in equilibrium to enjoy approximately the utility that she could get by interacting with a single perfectly aligned model Bob, without any additional assumptions on how close that utility is to her first best. To do this, we modify the design of the game.

### v2

@@ Abstract
We complement the theory with two forms of empirical evidence: First, we test our alignment condition on both synthetic and real-world data. We show that synthetically generated LLM utility functions (produced via perturbations of the same prompt to evaluate instances on a movie recommendation (MovieLens) and ethical judgement (ETHICS) dataset) quickly produce a convex hull that contains a good approximation of a given utility function even when none of the individual LLM utility functions is well aligned. We show similar findings using human and LLM responses on real-world polling data (OpinionQA): a convex hull of LLM opinions can approximate human opinions more accurately than any individual LLM across a wide range of survey questions. Second, we perform simulations of the best-AI selection game using best response dynamics, which show that competition among individually misaligned agents reliably improves user utility when the approximate convex hull assumption is satisfied.
@@ 1 Introduction
Aligning a single AI model to the objectives of its user is a hard problem, not just because of technical complexity, but because the incentives of AI designers may themselves be misaligned with users. But does our inability to solve the alignment problem preclude our ability to get the benefits of interacting with a strong aligned model? In this paper we study a setting in which it does not: when we may interact with multiple differently misaligned models in a strategic setting. In particular, we study settings in which there are many AI models available. They are produced by providers like e.g. OpenAI, Anthropic, Google, Meta, AWS, and xAI. These companies produce models reflective of their own incentives, none of which are necessarily well aligned to their user. We note that there has already been significant concern that the designers of LLMs are training them to influence users towards the politics of their creators (Menn, 2025; Kay, 2025; Gilbert, 2024; Hackenburg et al., 2025). In lieu of alignment of any model, we assume instead a much weaker condition on the entire market of providers: that (for a well specified task) an approximation of the user’s utility function lies somewhere in the convex hull of the utility functions of each of the AI companies. This is a condition that does not require that any single model is optimizing a utility function that is similar to that of the human user, and becomes a weaker assumption the more differently aligned models there are that are available to use. We remark at the outset that we primarily use the language of alignment of the human designers, and speak as if these designers are the strategic actors — but we could also think about the agents training and developing these AIs as themselves being AIs, whose individual misalignment results from the difficulty of the technical alignment problem. Having AI models themselves involved in the AI training process is a prominent part of thinking about the development of “super-intelligence” (see e.g. (Kokotajlo et al., 2025)) and is already part of current practice in limited ways (Leike et al., 2018; Bai et al., 2022).
@@
There are many ways that our approximate market alignment assumption could arise amongst competing AI providers. Consider a near-future scenario in which a human doctor has access to predictive medicine LLMs able to aid in the diagnoses and treatment of patients. The goal of the human doctor might be to provide the best treatment possible for her patients. The LLMs on the other hand might opt for better treatments all else being equal, but might also prefer to prescribe drugs manufactured by a particular company (say if that drug company is the creator or financial sponsor of the model). This preference results in a significantly misaligned model. However, since each drug has a single manufacturer, the “misaligned portion” of the AI model utilities is zero sum, and if all of the relevant drug companies participate in the predictive medicine LLM market, the doctor’s utility function will be in the convex hull (in fact the simple average) of the AI model utilities.
@@
We study how, in settings where approximate market alignment holds, strategic interactions between different models or model providers can allow the human user to realize the full benefit of interacting with a single perfectly aligned model by interacting with many differently misaligned models. While most AI safety research focuses on aligning individual systems or cooperative multi-agent approaches, we study how the benefits of perfect alignment can emerge from market-like competition among misaligned AI providers.
@@ 1.1 Our Model and Results
    In Section 6 we conduct a simple stylized experiment designed to test our core premise that given a set of AI models, there may be a utility function in the convex hull of the set of all AI agent utility functions that is substantially better aligned than any of the individual AI utility functions themselves. We test this premise in experiments on a few datasets. In the first we simulate a “human” utility function by using an LLM with a hand-crafted prompt, and ask it to evaluate 1000 ethical scenarios from the ETHICS dataset (Hendrycks et al., 2021). To simulate “AIs” that are designed to be aligned with the human utility function but are only noisy approximations, we produce perturbations of the original (“human”) prompt by asking a language model to rephrase the prompt while maintaining its core intent. We produce 100 such perturbations, resulting in up to 100 “AI personas” that we also use to evaluate the same 1000 ethical scenarios. Finally as a function of the number of AI models $k$ we evaluate the alignment of 1) the best aligned of the $k$ AI personas and 2) the best aligned utility function that can be computed within the convex hull of the $k$ AI personas. We repeat the experiment on the MovieLens dataset (Harper and Konstan, 2015) in which we use the average human annotation of movies as the “human” utility and similarly simulate 100 AI utility functions through 100 variations of a prompt. Finally, we move to real human preference data, using an existing human-LLM polling dataset of public opinion survey responses (Santurkar et al., 2023). Here, we treat individual human responses as human utilities and LLM responses as AI utilities. On all datasets we find that the best utility function in the convex hull of the AI utility functions is substantially better aligned to the “human” than any of the AI personas themselves. This supports our main conceptual contention that the target of alignment within the convex hull of many models may be substantially easier to obtain than alignment of any single model individually. Finally, we conduct a simple stylized experimental evaluation of the equilibria of a variant of the game we study in Section 5. We compute equilibria of a tractable special case of this game with a variety of utility functions, and find that consistent with our theory, Alice’s utility in equilibrium is always at least as high as predicted by our theory based on the alignment error of the best approximation to her utility function in the convex hull of the AI providers utilities — and sometimes substantially better.
@@ 1.2 Related Work
Bayesian Persuasion was introduced by Kamenica and Gentzkow (2011) — in the canonical model, there is a single informed “sender” and an uninformed “receiver” who share a common prior. The sender commits to a “signaling scheme”, which is a mapping from observations to messages sent to the receiver, who conditions on the message and takes their best response action under their posterior. We adopt the basics of this model, but extend it by allowing that both parties be differently informed, and that interaction involve a multi-round conversation rather than a single message. Multi-sender Bayesian Persuasion was introduced by Gentzkow and Kamenica (2016) and studies the standard Bayesian Persuasion model with multiple senders who simultaneously commit to a signaling scheme (playing, as in our paper, a simultaneous move commitment game). A number of papers have since studied multi-sender Bayesian Persuasion (Gentzkow and Kamenica, 2017; Li and Norman, 2018; Au and Kawai, 2020; Wu, 2023). We focus here on the most relevant papers.
@@
Ravindran and Cui (2020) study competing senders with zero-sum preferences over a receiver’s beliefs. They show that competition leads to full revelation of the state in all equilibria, provided the senders’ utility functions are “globally nonlinear”. This technical condition can hold in a standard receiver model only if the receiver has a different optimal action for every distinct state of the world. This condition cannot hold whenever e.g. the number of states of the world exceeds the number of actions. Our work does not assume that the leaders Bob are engaged in a zero-sum game with each other — rather our market alignment assumption can be viewed as assuming that the misaligned portions of their utility functions are approximately zero-sum under some non-negative reweighting. We also do not require an analogue of the “globally nonlinear” assumption, and so our results can apply to settings with large state spaces.
@@
Gradwohl et al. (2022) study a Bayesian persuasion game in which a receiver chooses to interact with only one of several competing senders (similar to our model in Section 5). As we do, they find that competition can force senders to be fully informative in equilibrium. In addition to the greater generality of our setup beyond Bayesian persuasion, our work differs in its core assumptions. The assumption driving the results of Gradwohl et al. (2022) is that the senders are uncertain about each other’s utility functions, and that any sender has a non-zero probability of being perfectly aligned with the receiver. We instead introduce and use the arguably more general “approximate market alignment” assumption, which only requires the user’s utility to lie within the convex hull of the AI agents’ utilities — we do not require any uncertainty about the AI agent utility functions, or any possibility that any of them are individually aligned with the user. Hossain et al. (2024) study the problem of multi-sender Bayesian Persuasion from a computational perspective, and prove worst-case hardness results for both the receiver’s best-response problem and for the senders’ equilibrium computation problem. They also design and evaluate neural network architectures suited to the (heuristic) computation of equilibria in such games.
@@ 2 Preliminaries
This section establishes the formal framework for our analysis. We first introduce the players and their information structure, then present our key modeling assumption about approximate market alignment, and finally define the communication protocol and game structure.
@@ 2.1 Players and Information Structure
2.1.1 The Market Alignment Assumption
We now turn to our key modeling assumption: that Alice’s utility can be approximately represented as a weighted combination of the AIs’ utilities, a notion we call “market alignment.”
Definition 1 (Approximate Market Alignment).
@@ 2.1.1 The Market Alignment Assumption
This assumption is central to our results.
@@ 2.2 Communication Protocol and Game Structure
With this in place, we can now define the communication protocol that governs how Alice and the AIs interact.
@@ 2.2.2 The Communication Protocol
The communication protocol models realistic constraints on human-AI interaction: conversations have limited rounds, messages have bounded complexity, and the human must process information from multiple AIs simultaneously. Alice engages in parallel private conversations with each AI, which captures settings where she can query multiple models independently. In most of the paper, Alice engages in $R$ rounds of private, parallel conversations with each of the $k$ Bobs (we will change the protocol in Section 5). Let $M$ be the message space.
@@
Conversation Rule Examples. For the Bobs, conversation rules correspond to the deployed model policy: for example, “Model X with provider Y’s safety layer and refusal policy.” Alice observes this conversation rule indirectly by knowing which API or product she is using, and how it has responded in the past to her and to others.
For Alice, a conversation rule is her repeated interaction strategy with the models. This interaction strategy may be adaptive to new information, not only within a conversation with one model, but across conversations. Examples include:
  • “Ask all models the same question.”
  • “Ask all models the same question, then pick the model with the promising answer and ask adaptive follow-up questions only to that model.”
  • “Ask all models the same question, then show the most promising answer to the other models and ask for critiques.”
  • “Phrase the question in $k$ different ways and ask each of the models a differently-phrased question.”
@@ 3.1 The Identical Induced Distribution Condition
Here, $\vec{C_{B}^{-i}}$ denotes the vector of all other Bobs’ strategies, and $\mathcal{I}^{*}((\vec{C_{B}^{-i}},C_{B}^{*}))$ is the induced distribution when Bob $i$ unilaterally deviates to $C_{B}^{*}$.
@@
Note that Proposition 1 is a joint condition on Alice’s utility and the expressiveness of the Bobs’ message spaces, and is fully independent of the Bobs’ utilities.
@@ 3.2 Strategic Competition Leads to Near-Optimal Outcomes
We can now state our first result: under the identical induced distribution assumption, approximate market alignment implies that in equilibrium, Alice gets utility approximately that of interacting with a single, perfectly aligned leader. In particular, if the message space is expressive enough to allow an aligned leader to communicate to Alice her Bayes-optimal action, approximate market alignment is sufficient for Alice to obtain approximately her first-best utility.
@@
If the multi-leader game satisfies the identical induced distribution condition, and if the Bobs satisfy the $\varepsilon$-market alignment condition, then Alice’s expected utility in any Nash equilibrium is at least $U_{A}(C_{B}^{*})-2\varepsilon$.
@@
Now we use the approximate market alignment assumption, which states that $\sum w_{i}U_{i}(a,y)$ is $\varepsilon$-close to $u_{A}(a,y)-c$. For the left-hand side:
@@ 4.2 Equilibrium Analysis Under the $(\delta,C_{B}^{*})$-Close Condition
Suppose the leaders Bob satisfy the $\varepsilon$-market alignment condition and the game satisfies the ($\delta,C_{B}^{*}$)-close condition (Definition 13) for a reference strategy $C_{B}^{*}$. Let $U_{A}(C_{B}^{*})$ be Alice’s expected utility from interacting with a single Bob using $C_{B}^{*}$. Then in any Quantal Response Nash Equilibrium (Definition 12), her expected utility is at least:
@@
The right-hand side can be bounded using the market alignment assumption:
@@ 4.3 From Information Substitutes to Utility Guarantees
Finally we are in a position to put all of the pieces together. If Alice is non-strategic (in that she commits to using the straightforward conversation rule, and the quantal response decision rule), and if in addition the underlying distribution satisfies the information substitutes condition, then if the Bobs satisfy market alignment, then Alice obtains close to her first best utility in every Nash equilibrium.
@@
Suppose the underlying distribution satisfies the Information Substitutes Condition (Definition 14) and the leaders Bob have an average market alignment error of $\varepsilon$. If Alice commits to the straightforward conversation rule and a $\lambda$-quantal response decision rule, her expected utility in any Quantal Response Nash Equilibrium of the induced game is close to the first-best optimal utility:
@@
Here, $\varepsilon$ is the alignment error from the market alignment assumption (Definition 1).
@@ 5 Winning the User: Assumption Free Guarantees
In Section 3 we showed that Alice could obtain her first-best utility in equilibrium amongst AI models Bob who satisfy the market alignment assumption, assuming that a single perfectly aligned Bob could cause Alice to enjoy her first best utility. In Section 4, we showed that if Alice is non-strategic and uses quantal response rather than best response, then the assumption that a perfectly aligned Bob could cause Alice to enjoy her first best utility could be relaxed to an approximate version. In this section, we give a setting in which Alice is guaranteed in equilibrium to enjoy approximately the utility that she could get by interacting with a single perfectly aligned model Bob, without any additional assumptions on how close that utility is to her first best. To do this, we modify the design of the game.